------------------------------------------------------------
Date: Sat Aug 26 15:56:34 2000, Shift: 08:00-16:00
Supervisor: R. Debbe, with: RD BA
Log book#5 pages 20 - 25
Shift Summary:
------------------------------------------------------------
1. Machine Status:  Injection started at 8:50 and beam cogged
at 9:10 Good intensities (Blue:14.3x10^9/Yellow:20x10^9)
ZDC scaler 1 counting at 300 Hz.
Currents in TPCs are high; TPM2 and T2A reach 2 micro Amps
 
At 11:05 MCR announced that the beam will be dumped because
of "cryo. problems"
Since then, not a hint of beam. The message from MCR says
that they have a "Yellow quench link interlock with no estimate"
2. Achievements: 
Polarities of D1 and D2 were flipped from B to A before taking
data.
Wrote runs 2445 (200K trig 1) MRS 45 deg 4Kgauss pola A
           2446 (144K trig 1) same as above
    
run 2447 is a TPC pedestal run.
3. No detector problems. The currents in TPCs go back to ~0
when beam is off.
Some SuperMon plots for TOFW H1 and H2 are empty.
D4 was turned on for 3 hours to test it.
------------------------------------------------------------
This archive was generated by hypermail 2b29 : Sat Aug 26 2000 - 15:59:55 EDT